### v1

# Omni2Sound: Towards Unified Video-Text-to-Audio Generation

Abstract
footnotetext: Corresponding author.
[…]
1 Introduction
Early audio generation models typically rely on unimodal conditioning. Text-to-Audio (T2A) [kreuk2022audiogen, liu2023audioldm, evans2024stableaudioopen, ghosal2023instruction] offers strong semantic fidelity and generalization but lacks dense temporal control. Conversely, Video-to-Audio (V2A) [luo2023diff-foley, wang2024frieren, zhang2024foleycraft, chen2024videoguided] ensures fine-grained temporal synchronization with video, yet suffers from weak reasoning in complex scenes and unfaithful generation (e.g., unexpected music or speech) [bain2025vintage, yu2024reasonaudio]. To address this, recent Video-Text-to-Audio (VT2A) methods [chen2024hunyuanfoley, liu2025thinksound, cheng2025mmaudio, tian2025audiox, bain2025vintage] jointly condition on video and text. While VT2A achieves strong both semantic understanding and temporal alignment, its reliance on simultaneous inputs constrains its applicability [tian2025audiox]. Crucially, most VT2A systems lack robustness [liu2025thinksound, chen2024hunyuanfoley, bain2025vintage], degrading sharply under missing-modality conditions (video-only or text-only).
These constraints motivate a unified framework natively supporting VT2A, V2A, and T2A within a single model. This unified paradigm aligns with the AIGC shift, eliminating the redundant architectures and deployment complexity of hard-switching between specialized models. Recent work has begun to advance this unified approach. MMAudio [cheng2025mmaudio] introduces a multimodal joint training framework to improve V2A generation, optionally conditioning on text using large-scale text–audio pairs. Moreover, AudioX [tian2025audiox] enhanced flexibility by supporting broader modality combinations. Despite this progress, two challenges in the unified VT2A framework remain underexplored.  [^1] https://swapforward.github.io/Omni2Sound
Figure 1: Challenges in scaling high-quality audio captions.
First, there is a scarcity of high-quality audio captions that are well-aligned with both audio and video cues. Most unified or specialized VT2A studies create their (V, T, A) training triplets by pairing videos (V) and their audio (A) with captions (T) generated solely from the audio [tian2025audiox, chen2024hunyuanfoley]. However, this approach introduces severe semantic conflict in the multimodal training data (see Figure 1): a frequent mismatch between the visual content and the (audio-only) text caption. This conflict is rooted in the audio modality’s inherent ambiguity (e.g., a tennis hit vs. distant fireworks, or car engine noise vs. an electric drill). This fundamental ambiguity is then exacerbated by the limited capabilities of earlier audio-language models, which are prone to severe hallucinations (e.g., omissions and mislabels) [chen2025detecting]. In our preliminary experiments, we found these modality conflicts caused by mismatches between V-T conditions directly lead to unstable convergence and a significant degradation in audio faithfulness. Unfortunately, there is still a lack of high-quality V-T-A triples for unified VT2A models training, as we further discuss in Section 2.
Second, two critical types of task competition within unified VT2A frameworks remain underexplored. (1) Cross-Task Competition. Prior work, notably MMAudio [cheng2025mmaudio], established that incorporating T-A pairs enhances the generalization and quality of V2A generation. However, training a unified model to excel at both V2A and T2A presents a significant challenge: as shown in our preliminary experiment (Table 5), this joint training introduces a severe T2A-V2A adverse trade-off, rooted in the heterogeneity between text and video modalities. Prioritizing one task during training consistently degrades the performance of the other, indicating a zero-sum optimization dynamic. (2) Intra-Task Competition. We also observe competition within the VT2A task itself. This competition manifests as a modality bias during generation process that undermines cross-conditional consistency, revealing two key failure modes: a bias towards text leads to poor A-V synchronization (Table 6), while a bias towards video exhibits low text-audio faithfulness in off-screen generation scenarios (Table 7).
To address data scarcity, we first introduce SoundAtlas in Section 3, a large-scale, agent-generated multimodal audio-caption dataset. It augments the two largest audio datasets, VGGSound [chen2020vggsound] and AudioSet [gemmeke2017audio], providing semantically rich and temporally detailed captions that even surpass human-expert quality (Table 2). Built on current advanced multimodal foundation models (Gemini-2.5 Pro [team2023gemini] and Qwen-2.5-VL [qwen2025vl]), we develop a multi-turn, agentic annotation pipeline featuring a junior–senior agent handoff, vision-to-language compression, and post-hoc hallucination filtering. This pipeline delivers cost-controlled annotations while maintaining tight visual–audio–text (V–A–T) alignment and a markedly higher text-audio faithfulness than prior datasets. Interestingly, we find its quality is high enough to even correct human annotation errors in VGGSounder [zverev2025vggsounder].
Building on this dataset, we propose Omni2Sound in Section 4, a diffusion-based unified model supporting flexible input modalities while maintaining both audio-visual synchronization and high-fidelity generation. To address cross-task and intra-task competition, we introduce a three-stage progressive training schedule that departs from naive joint training. First, a large-scale T2A pretraining stage establishes a robust generative prior, enabling minimal high-quality T2A replay in the subsequent stage to prevent catastrophic forgetting. Subsequently, our Multi-task Interleaved Training integrates V2A and T2A tasks with high-quality VT2A triplets. Our central insight is that this VT2A data serves as a semantic bridge: by aligning the heterogeneous feature spaces of video and text, it effectively converts zero-sum cross-task competition into a cooperative optimization dynamic, thereby mitigating training resource contention. To resolve the intra-task competition, our third stage employs a decoupled Robustness Training. We utilize two synergistic augmentations to balance cross-modal reliance: Text Dropout penalizes text bias to enhance A-V synchronization, while Off-screen Synthesis counteracts video bias to ensure textual faithfulness. This decoupled approach rectifies key failure modes, maintaining high-fidelity generation even in challenging, asymmetric input scenarios.
Figure 2: Data Construction Pipeline of SoundAtlas (Left). Comparison against SOTA baselines and human annotations (Right) .
Finally, we construct VGGSound-Omni in Section 5, the first comprehensive benchmark to establish a unified evaluation standard for VT2A, V2A, and T2A. It provides high-quality, human-verified annotations for all three tasks and introduces a challenging off-screen audio generation track. As a result, with a vanilla DiT [Peebles2023DiT] backbone, Omni2Sound achieves unified state-of-the-art performance across all three tasks against both unified and specialized models, showing high-fidelity audio quality, tight audio-visual synchronization, and excellent generation faithfulness.
2 Related Works
Audio Caption Dataset.
Human-annotated benchmarks like AudioCaps [kim2019audiocaps] (46k) and Clotho [drossos2020clotho] (5k) offer high-quality alignment, but their limited scale, high cost, and lack of detail make them unsuitable for training modern, large-scale models. Automated pipelines emerged to address data scarcity. WavCaps [mei2023wavcaps] used LLMs to refine noisy web metadata (400k captions), and AudioSetCaps [lipping2024audiocaps] used ALMs+LLMs to extract and aggregate details from audio, speech, and music, significantly increasing data volume. As detailed in our Introduction, these audio-only methods suffer from high hallucination rates and can lead to cross-modal conflicts that destabilize VT2A training, stemming from the audio modality’s inherent ambiguity. Visually-enhanced (VE) annotation pipelines like Auto-ACD [han2024autoacd] and Sound-VECaps [yu2024soundvecaps] emerged to leverage visual cues for cross-modal constraint. While promising, existing implementations adopt a separate-then-fuse pipeline: unimodal models extract separate textual cues (e.g., image captions, audio tags), which are then merged by a final LLM. This pipeline is suboptimal, as the LLM fuses lossy textual representations, not raw modalities, leading to the accumulation and amplification of unimodal hallucinations. While using native end-to-end multimodal models (e.g., Gemini [team2023gemini]) seems a natural solution, it also proves suboptimal. As we demonstrate in Section 3, this method faces prohibitive costs and a pervasive visual bias that prevents truly audio-centric captioning. There remains a lack of a large-scale, high-quality visual–audio–text (V–A–T) aligned audio caption dataset suitable for training unified VT2A models.
Unified Audio Generation Model.
The audio generation paradigm is shifting towards unified, omni-modal frameworks, a trajectory initiated by MMAudio [cheng2025mmaudio]. While it integrated V2A and T2A, its approach was fundamentally V2A-centric, using T-A pairs merely as augmentation for V2A rather than optimizing T2A as a co-equal task. Subsequent works like AudioX [tian2025audiox] and AudioGen-Omni [wang2025audiogenomni] expanded this scope to more flexible modality combinations. However, these efforts often relied on brute-force data scaling (e.g., AudioX with over 9 million samples), which revealed inefficiencies and failed to yield proportional SOTA performance. Critically, these early models [cheng2025mmaudio, tian2025audiox, wang2025audiogenomni] largely overlooked the inherent cross-task competition stemming from co-training these diverse sub-tasks. UniFlow-Audio [xu2025uniflow] is the first to systematically address this by categorizing tasks into Time-Aligned (TA) and Non-Time-Aligned (NTA) classes and analyzing their competitive dynamics. However, its analysis remains coarse-grained, failing to investigate the granular competition within the TA category (i.e., V2A vs. T2A). Moreover, the challenging case of joint cross-modal generation (VT2A) remains unaddressed. Consequently, a fundamental study on task competitive dynamics within a unified VT2A framework remains absent.
3 SoundAtlas: V-A-T Data Construction
Existing automated audio caption datasets often suffer from severe visual-audio-text (V-A-T) misalignment with high hallucination rates due to the limitations of early ALMs [lipping2024audiocaps, han2024autoacd, yu2024soundvecaps]. While recent native multimodal foundation models like Gemini 2.5 [team2023gemini, MMAR:2025, Qwen3Omni:2025] offer strong capabilities, we find that a naive implementation—processing raw video-audio pairs directly—is suboptimal for audio caption dataset construction. Specifically, it incurs prohibitive costs (approx. \$10,275 per 1M samples; see Appendix A) and suffers from inherent visual bias, where models hallucinate auditory labels for non-existent events due to visual interference, as shown in Figure 1.
To address these challenges, we introduce SoundAtlas, constructed via a cost-effective, multi-turn agentic annotation pipeline. As illustrated in Figure 2, our pipeline integrates vision-to-language compression to mitigate visual bias, a junior–senior agent handoff to optimize cost-efficiency, and rigorous post-hoc filtering to ensure annotation fidelity. Full prompt instructions are detailed in Appendix B.
A-V Consistency Routing.
We first apply A-V Consistency Routing on raw video from AudioSet [gemmeke2017audio] and VGGSound [chen2020vggsound]. This step is based on the core finding that visual cues are reliable for high-consistency A-V clips but act as distractors in low-consistency clips as shown in Figure 2. We classify samples based on ImageBind alignment ($s_{ib}$) using thresholds $\tau_{\text{low}}=0.20$ and $\tau_{\text{high}}=0.30$: (i) High-consistency ($s_{ib}>\tau_{\text{high}}$) enter the A-V Enhanced Path; (ii) Medium-consistency ($\tau_{\text{low}}\leq s_{ib}\leq\tau_{\text{high}}$) are routed to the Audio-Only Path to prevent visual hallucinations; and (iii) Noise ($s_{ib}<\tau_{\text{low}}$) is discarded.
Table 1: Semantic Faithfulness (CLAP Score) of Different Data Construct Pipelines on AudioSet and VGGSound.
Method AudioSet VGGSound
LA-CLAP $\uparrow$ MS-CLAP $\uparrow$ LA-CLAP $\uparrow$ MS-CLAP $\uparrow$
AudioSetCaps [lipping2024audiocaps] 0.330 0.397 0.351 0.421
Sound-VECaps [yu2024soundvecaps] 0.370 0.425 - -
Auto-ACD [han2024autoacd] 0.396 0.437 0.409 0.457
SoundAtlas (Ours) 0.447 0.485 0.461 0.502
Table 2: Caption quality comparison via MLLM-as-a-judge and human evaluation, reporting the Mean Win Rate for Semantic (MWR-S) and Temporal (MWR-T) alignment. Human-Expert refers to the human-annotated captions from AudioCaps [kim2019audiocaps].
Method MLLM Evaluation Human Evaluation
MWR-S $\uparrow$ MWR-T $\uparrow$ MWR-S $\uparrow$ MWR-T $\uparrow$
Auto-ACD [han2024autoacd] 0.39 0.41 0.31 0.26
Human-Expert [kim2019audiocaps] 0.36 0.51 0.46 0.55
SoundAtlas (Ours) 0.75 0.58 0.71 0.69
Vision-to-Language Compression.
This step implements our key insight: vision must be treated as a contextual constraint, not a primary input. We found that compressing the visual stream into a textual representation ($c_{v}$) is a more effective strategy, as it simultaneously addresses both of our defined challenges. First, it addresses cost by replacing the prohibitively expensive raw video input ($V+A$) with a cost-effective text-audio prompt ($c_{v}+A$). Second, it robustly mitigates cross-modal hallucinations by filtering the visual bias, providing only low-bias semantic context (e.g., "A man and a woman are standing…") rather than a misleading raw visual stream. Therefore, for samples $V$ routed to the A-V Enhanced Path, we use Qwen-2.5-VL [qwen2025vl] to analyze the video $V$ (without its audio $A$) and generate the textual representation $c_{v}=\text{Qwen}(V)$. Conversely, samples on the Audio-Only Path are assigned a null context.
Junior–Senior Agent Handoff.
All samples then enter our handoff pipeline. The task is first assigned to the Junior agent, $G_{\text{junior}}$ (Gemini 2.5 Flash), which receives the audio $A$ and the optional visual context $c_{v}$. Let the output caption be $c_{a}=G_{\text{junior}}(A,c_{v})$. This caption $c_{a}$ is then flagged if it (i) meets our complexity criteria (text-based heuristics to identify complex audio scenes), (ii) contains high-frequency hallucination phrases, or (iii) fails our differentiated CLAP [elizalde2023clap] check, $\text{CLAP}(c_{a},A)<\tau_{clap}$, where $\tau_{clap}$ is $0.35$ for general audio and $0.15$ for music. Flagged tasks are escalated to the Senior agent, $G_{\text{senior}}$ (Gemini 2.5 Pro). To control costs, the Senior agent’s reasoning output is limited to 128 tokens, providing a more precise caption.
Post-hoc Filtering and Verification.
Finally, all generated captions $c_{a}$ undergo a two-stage verification. First, a CLAP (T-A) filtering model [elizalde2023clap] ensures high Text-Audio faithfulness; captions where $\text{CLAP}(c_{a},A)<\tau_{verify}$ are discarded. Second, for captions from the A-V Enhanced Path ($c_{v}\neq\emptyset$), an A-V-T Verifier, $V_{\text{AVT}}$, ensures $c_{a}$ is a reasonable acoustic inference given $c_{v}$. Captions that pass all filters are accepted into the final dataset $\mathcal{D}_{\text{SoundAtlas (Ours)}}$, which augments VGGSound [chen2020vggsound] and AudioSet [gemmeke2017audio] datasets with human-expert-level audio captions.
3.1 Comparison with Existing Pipeline
We compare SoundAtlas against other automated pipelines [lipping2024audiocaps, han2024autoacd, yu2024soundvecaps] on high audio-visual consistency subsets sourced from AudioSet and VGGSound, where ImageBind score $s_{ib}>0.30$. As shown in Table 1, SoundAtlas significantly outperforms all competitors on both LA-CLAP and MS-CLAP scores, demonstrating superior text-audio alignment. Additionally, we conduct a fine-grained MLLM-as-a-judge (Gemini 3.0 Pro [team2023gemini]) evaluation on the intersection of AudioCaps and all compared datasets [kim2019audiocaps]. As shown in Table 2, SoundAtlas achieves a substantially higher mean win rate in semantic alignment (MWR-S) and temporal alignment (MWR-T) than both the strongest baseline (Auto-ACD) and the Human-Expert annotations, across both semantic and temporal alignment. To mitigate potential evaluation bias, a follow-up human validation study is conducted, further corroborating our results (details in Appendix Section C). As illustrated in Figure 2 (right), SoundAtlas demonstrates clear superiority over existing automated datasets, characterized by its richer semantic content and explicit temporal ordering.
4 Omni2Sound: Unified VT2A Generation
Building on SoundAtlas, we propose Omni2Sound, a Diffusion-based unified VT2A model supporting collaborative (VT2A) and unimodal (V2A, T2A) control.
4.1 Foundation Model Architecture
We adhere to a principle of simplicity and scalability, adopting a standard Diffusion Transformer (DiT) backbone [Peebles2023DiT] conditioned on latent features from a pre-trained audio VAE [evans2024fast]. As shown in Figure 3, the backbone is conditioned on multimodal inputs using a decoupled injection approach, which is separated into two distinct branches: (1) Semantic Branch (What) and (2) Temporal Branch (When). To capture global semantic context, we concatenate text embeddings from Flan-T5 [chung2024scaling] ($F_{t}$) and visual features from CLIP [radford2021learning] ($F_{v}$, sampled at 8 fps) along the temporal dimension, which are then injected via cross-attention layers. Crucially, this design allows for flexible unimodal generation (V2A or T2A) by simply omitting the absent modality without requiring padding constraints. For the Temporal Branch, to ensure fine-grained synchronization, we follow [cheng2025mmaudio] to utilize a Synchformer [Viertola2024Synchformer] to extract dense visual-temporal features ($F_{s}$) and then inject it globally via Adaptive Layer Normalization (AdaLN).
This decoupled architecture effectively (1) achieves the flexibility of multi-condition frameworks like AudioX [cheng2025mmaudio], supporting extensible conditions without architectural modification; and (2) maintains precise temporal alignment comparable to MMAudio [cheng2025mmaudio] (powered by its well-designed MM-DiT architecture).
4.2 Three-stage Progressive Multi-task Training
As established in Section 1, native joint training faces Cross-Task and Intra-Task competition. To resolve both, we design the following three-stage progressive schedule. To resolve both, we design a three-stage progressive multi-task training schedule.
Stage 1: Large-scale T2A Pretraining.
We first conduct standalone T2A pretraining on large-scale text-audio pairs without a quality filter. Following the latent diffusion framework [stableaudio2024, Peebles2023DiT], our DiT backbone (Section 4.1) is trained to progressively denoise a noisy latent $z_{t}$ at timestep $t$, conditioned on text embeddings $H_{c}$. The model, $\epsilon_{\theta}$, is optimized via the simple L2 loss:
$L=\mathbb{E}_{t,z_{t},\epsilon}\|\epsilon-\epsilon_{\theta}(z_{t},t,H_{c})\|^{2}$
Figure 3: Overview of our unified VT2A framework, which integrates global semantics and temporal alignment, supporting flexible T2A, V2A, and VT2A generation.
This pretraining provides two benefits: first, it establishes a robust generative prior before introducing the heterogeneity of video conditions; second, it allows for significantly reduced T2A sampling frequency in the next stage without suffering catastrophic forgetting, thereby mitigating resource contention.
Stage 2: Multi-task Interleaved Training.
This stage resolves Cross-Task Competition using a Multi-task Interleaved strategy with Task-Balanced Sampling. At each step, a single task $s\in\{V2A,T2A,VT2A\}$ is sampled from a categorical distribution $\text{Cat}(\pi)$, and a minibatch is drawn exclusively from its dataset $D_{s}$ for a single-task gradient update. This approach stabilizes optimization by avoiding within-batch loss mixing. This strategy is grounded in two key findings, which we validate experimentally (Section 6.3): (i) As demonstrated in our ablation study (Table 5), we find the VT2A task acts as a critical bridge. Adding it mitigates the adverse V2A-T2A trade-off, enabling their simultaneous optimization rather than a zero-sum competition. (ii) Supported by this bridge, we also found (Table 5) that a low sampling frequency (e.g., $\pi_{T2A}=0.1$) of high-quality T2A data is sufficient to prevent catastrophic forgetting. These findings allow our Stage 2 schedule to be driven primarily by video-conditioned tasks (V2A and VT2A), using T2A only minimally to retain its strong generative prior.
Stage 3: Intra-Task Resolution via Robustness Training.
While Stage 2 resolves the overarching Cross-Task Competition, the inherent Intra-Task Competition (modality bias) persists, particularly in challenging scenarios like off-screen generation. We therefore introduce a final, decoupled Robustness Training stage. This decoupling is essential: as we empirically demonstrate in Table 6, introducing robustness augmentations prematurely into Stage 2 destabilizes the fragile optimization process. Our decoupled approach, in contrast, is strategically designed to enhance cross-modal consistency without compromising the generative quality already achieved.
This stage employs complementary augmentations to create a balanced reliance on both modalities: (i) Text Dropout. By randomly deleting tokens from the text prompt, we create ambiguity that compels the model to rely more on the visual stream; strengthens A-V synchronization by counteracting a bias towards text. (ii) Off-screen Synthesis. Mixing in off-screen audio and augmenting the text prompt to describe it, we create samples where the audio is not represented by the video. This forces the model to rely more on the text condition, improving textual faithfulness against a video bias in off-screen audio generation.
Table 3: Comparison on VGGSound-Omni benchmark: Omni2Sound against SOTA models on T2A, V2A, and VT2A tasks. The w/ Video-LLaMA caps row evaluates Omni2Sound’s generalization to unseen captions generated by Video-LLaMA [zhang2023video-llama].
Task Method Distribution Matching Audio Quality Modality Alignment
KL FD FAD $\mathrm{\textbf{FD}_{\textbf{PaSST}}}$ PQ IS DS IB MS-CLAP
T2A AudioX [tian2025audiox] 1.68 9.04 1.42 109.94 6.37 15.15 - - 0.49
MMAudio [cheng2025mmaudio] 1.92 8.62 1.63 101.66 5.84 14.30 - - 0.50
[…]
    w/ Video-LLaMA caps 1.60 6.92 1.23 83.91 6.38 16.01 - - 0.51
V2A V-AURA [viertola2024vaura] 2.28 16.43 2.34 245.25 5.74 10.82 0.69 0.28 0.32
Frieren [wang2024frieren] 2.73 12.13 1.23 123.75 5.82 11.32 0.86 0.21 0.31
AudioX [tian2025audiox] 2.96 12.73 1.42 121.82 6.17 13.34 1.22 0.24 0.34
MMAudio [cheng2025mmaudio] 2.11 5.65 0.81 69.33 5.72 11.85 0.48 0.28 0.43
[…]
VT2A ThinkSound ($w/o$. CoT) [liu2025thinksound] 1.60 7.41 1.10 116.08 6.21 11.73 0.53 0.26 0.43
HunyuanVideo-Foley [chen2024hunyuanfoley] 1.74 10.02 2.36 100.53 6.18 11.58 0.57 0.32 0.45
AudioX [tian2025audiox] 1.59 8.29 1.24 103.37 6.17 14.94 1.23 0.26 0.49
MMAudio [cheng2025mmaudio] 1.63 5.28 0.91 68.44 5.84 13.44 0.49 0.29 0.49
[…]
    w/ Video-LLaMA caps 1.56 3.37 0.66 53.73 6.11 15.74 0.50 0.34 0.49
5 VGGSound-Omni: Unified Evaluation
A significant challenge in evaluating unified Video-Text-to-Audio (VT2A) models is the absence of a comprehensive benchmark. The VGGSound test set [chen2020vggsound] only provides sparse event labels and lacks detailed captions. Although recent work like VGGSounder [zverev2025vggsounder] significantly improved this by correcting and adding crucial modality labels (e.g., A, V, AV) for fidelity evaluation, it still lacks human-expert-level captions. To address this gap, we construct VGGSound-Omni, a new multi-track benchmark derived from the original VGGSound test set, designed for both standard unified and specialized off-screen VT2A tasks evaluations. The construction process is detailed below.
VGGSound-Omni Construction.
Our first step was to establish a high-fidelity, human-level caption set for all 14,000 videos, forming the primary evaluation track. We first generated an initial caption using our agentic pipeline (Section 3). We then systematically validated this output via an AI-assisted verification workflow: GPT-5 [openai2025gpt5intro] was tasked to act as an auditor, checking if our captions semantically covered all the “A” and “AV” labels from VGGSounder [zverev2025vggsounder]. Samples flagged with a mismatch were routed for targeted human verification. During this manual audit process, we found most of these flagged discrepancies stemmed from annotation errors within the VGGSounder data itself (e.g., label redundancy and human annotation errors caused by visual interference). After manually correcting for these identified errors, we established our final, human-verified captions as the definitive ground truth (GT) for evaluating all three tasks (VT2A, V2A, and T2A).
Complementing the primary set, we construct a challenging off-screen track (1,048 items). We curated this subset from two sources: (i) Natural events, filtered from VGGSound for low A-V correspondence (via IB-Score [girdhar2023imagebind] and Desync-Score [cheng2025mmaudio]) while excluding background speech; and (ii) Synthetic music, formed by mixing aligned background clips from MusicCaps [agostinelli2023musiccaps]. More Details are provided in Appendix D.
Table 4: Comparison on the Kling-Audio-Eval: Omni2Sound against SOTA models on T2A, V2A, and VT2A tasks.
Task Method Distribution Matching Audio Quality Modality Alignment
KL FD FAD $\mathrm{\textbf{FD}_{\textbf{PaSST}}}$ PQ IS DS IB LA-CLAP
T2A AudioX [tian2025audiox] 2.73 19.43 3.32 171.60 5.98 12.15 - - 0.28
MMAudio [cheng2025mmaudio] 2.54 11.25 5.07 142.71 5.54 9.28 - - 0.28
[…]
V2A AudioX [tian2025audiox] 3.13 18.90 4.01 205.48 5.87 8.31 1.20 0.23 0.13
MMAudio [cheng2025mmaudio] 2.94 13.41 3.87 159.30 5.50 7.59 0.62 0.24 0.14
[…]
VT2A ThinkSound ($w/o$. CoT) [liu2025thinksound] 2.53 11.99 3.52 206.93 5.77 6.09 0.66 0.22 0.19
HunyuanVideo-Foley [chen2024hunyuanfoley] 2.13 8.06 3.58 94.64 6.04 8.17 0.55 0.34 0.23
AudioX [tian2025audiox] 2.39 14.26 3.16 149.37 5.97 10.23 1.21 0.23 0.26
MMAudio [cheng2025mmaudio] 2.41 10.12 4.90 129.21 5.53 7.46 0.59 0.25 0.20
[…]
6 Experiments
6.1 Experiment Settings
Datasets.
For T2A backbone pre-training, we use a large-scale corpus comprising the train set of audio datasets such as AudioCaps [kim2019audiocaps], WavCaps [mei2023wavcaps], Clotho [drossos2020clotho], AudioSet [gemmeke2017audio], VGGSound [chen2020vggsound], FSD50k [fonseca2022fsd50k], as well as music datasets including MSD [bertin2011million] and FMA [defferrard2016fma]. To maintain consistency, all audio is segmented into 10-second clips and resampled at 16 kHz. Following this, the model is trained for unified VT2A tasks using our proposed SoundAtlas (Section 5) and a high-quality, PQ-score-filtered T-A subset derived from the aforementioned pre-training corpus. More details of the implementation are provided in Appendix Section G. For evaluation, we compare Omni2Sound with SOTA models on three benchmarks: our proposed VGGSound-Omni (Section 5), Kling-Audio-Eval [wang2025audiogenomni] and AudioCaps test set [kim2019audiocaps]. We strictly ensure that these evaluation benchmarks are strictly disjoint from all data used in our training stages to prevent potential data leakage.
Evaluation Metrics.
We implement our objective evaluation using the standardized AV-benchmark toolkit [hkchengrex2024avbenchmark] on 8-second clips, following previous work [cheng2025mmaudio]. We assess quality across four critical dimensions [liu2023audioldm]. For Distribution Matching, we measure feature similarity between generated and ground-truth audio using Fréchet Distance (FAD [hershey2017cnn], $\mathrm{FD}_{\mathrm{PaSST}}$ [koutini2022passt], FD [kong2020panns]) and Kullback-Leibler divergence (KL, $\mathrm{KL}_{\mathrm{PaSST}}$). Audio Quality is assessed via Inception Scores (IS [salimans2016improved], $\mathrm{IS}_{\mathrm{PaSST}}$) and Production Quality (PQ [tjandra2025meta]) for aesthetics. Semantic Alignment evaluates text-audio consistency (CLAP [elizalde2023clap], MS-CLAP [wu2023large]) and video-audio alignment (IB [girdhar2023imagebind]). Finally, Temporal Alignment is measured using the Desynchronization Score (DS) predicted by Synchformer [iashin2024synchformer]. Detailed metric definitions and calculations are provided in the Appendix.
6.2 Main Results
Evaluation on VGGSound-Omni.
We present our main results on VGGSound-Omni benchmark in Table 3. To ensure a fair comparison, all baseline models are re-evaluated using their official checkpoints and the standardized AV-benchmark toolkit [hkchengrex2024avbenchmark], using the same video and text conditions. The results demonstrate that Omni2Sound achieves state-of-the-art performance across all three unified tasks (T2A, V2A, and VT2A) compared to both previous unified VT2A models (AudioX [tian2025audiox], MMAudio [cheng2025mmaudio]) and specialized models (e.g. ThinkSound [liu2025thinksound], HunyuanVideo-Foley [chen2024hunyuanfoley]). To further validate Omni2Sound’s generalization beyond our SoundAtlas captioning style, we evaluate it on the same VGGSound test clips but use the Video-LLaMA [zhang2023video-llama] captions from ThinkSound [liu2025thinksound]. As shown in Table 2 (w/ Video-LLaMA caps), while performance sees a slight degradation, our model’s scores still surpass all baselines, confirming its robustness to unseen captioning styles.
Generalization on Third-Party Benchmarks.
To validate generalization, we evaluate on Kling-Audio-Eval [wang2025audiogenomni] and AudioCaps [kim2019audiocaps] results in Table 4 and Appendix Table 7. On Kling-Audio-Eval, Omni2Sound remains highly competitive despite the domain gap (YouTube-sourced SoundAtlas vs. Kling’s professional video). While trailing HunyuanVideo-Foley [chen2024hunyuanfoley] in some metrics, which is expected given its massive data advantage (100k vs 2k hours), our model consistently outperforms other unified and specialized baselines across all tasks. Furthermore, on AudioCaps, Omni2Sound achieves top-tier performance against specialized T2A models, securing the best scores in distribution metrics ($\mathrm{KL}$, $\mathrm{FD}$) and semantic alignment ($\mathrm{CLAP}=0.36$), while remaining highly competitive in audio quality (PQ) and the $\mathrm{FAD}$ metric.
Subjective Evaluation.
To validate perceptual performance, we conduct a human evaluation (detailed in Appendix F) across three dimensions: Acoustic Fidelity (MOS-Q), Semantic Consistency (MOS-S), and Temporal Synchronization (MOS-T). As shown in Appendix Fig. 4, Omni2Sound outperforms all baselines on both VT2A and V2A tasks. Crucially, these subjective results are highly consistent with the objective metrics in Table 3, confirming our model’s superiority in both generation quality and cross-modal alignment.
Table 5: Ablation study on the Stage 2 multi-task training strategy. TA*/VTA* denotes data from our high-alignment SoundAtlas dataset, while TA/VTA denotes data from a baseline with audio-only captions generated by Gemini 2.5.
Training Strategy $\pi_{T2A}$ : $\pi_{V2A}$ : $\pi_{VTA}$ T2A Task V2A Task VT2A Task
FAD FD FAD FD DS IB FAD FD DS IB
TA+VA 0.20 : 0.80 : 0.00 1.36 5.52 0.56 4.13 0.50 0.33 - - - -
TA+VA 0.40 : 0.60 : 0.00 1.06 4.62 0.62 4.63 0.52 0.32 - - - -
TA*+VA+VTA* 0.10 : 0.35 : 0.55 0.94 4.22 0.57 3.61 0.49 0.33 0.53 2.83 0.51 0.32
TA+VA+VTA 0.20 : 0.30 : 0.50 1.13 4.68 0.56 4.22 0.50 0.32 0.62 3.51 0.51 0.33
Table 6: Ablation study on our progressive multi-task training. We compare our full S1 $\rightarrow$ S2 $\rightarrow$ S3 model against three baselines (S2, S1 $\rightarrow$ S2, and S1 $\rightarrow$ [S2+S3]). All models are trained for the same total 1.2M steps.
Task Method FAD FD DS IB
T2A S2 1.22 5.88 - -
S1 $\rightarrow$ S2 0.94 4.62 - -
S1 $\rightarrow$ [S2+S3] 1.11 4.45 - -
S1 $\rightarrow$ S2 $\rightarrow$ S3 1.01 4.61 - -
V2A S2 0.68 4.70 0.47 0.33
S1 $\rightarrow$ S2 0.57 3.61 0.49 0.33
S1 $\rightarrow$ [S2+S3] 0.60 3.81 0.47 0.34
S1 $\rightarrow$ S2 $\rightarrow$ S3 0.51 3.41 0.47 0.35
VT2A S2 0.63 4.40 0.49 0.33
S1 $\rightarrow$ S2 0.53 2.83 0.51 0.32
S1 $\rightarrow$ [S2+S3] 0.61 3.27 0.50 0.33
S1 $\rightarrow$ S2 $\rightarrow$ S3 0.53 2.95 0.49 0.34
6.3 Ablation Studies
We first analyze the multi-task training dynamics in Table 5 to demonstrate how high-quality data resolves task competition, and then use Table 6 to prove the necessity of our three-stage progressive training schedule.
High-Quality VT2A Data as a Critical Bridge.
We first investigate the Cross-Task Competition between V2A and T2A, which still persists even when models are based on the T2A pretraining from Stage 1. As shown in Table 5 (rows 1-2), a naive joint training of V2A and T2A results in a severe adverse trade-off. Increasing the T2A sampling ratio ($\pi_{T2A}$) from 0.20 to 0.40 improves T2A performance (FAD 1.36 $\rightarrow$ 1.06) but simultaneously degrades V2A generation (FAD 0.56 $\rightarrow$ 0.62), preventing simultaneous optimization.Our central insight is that this conflict is resolved by introducing high-quality VT2A data as a critical bridge. This hypothesis is validated in row 3, which introduces our SoundAtlas data (denoted by TA* and VTA*). The results show a dramatic performance boost, achieving the best metrics across all tasks (e.g., T2A FAD 0.94, V2A FD 3.61, VT2A FD 2.83). This confirms that the high A-V-T alignment in SoundAtlas is essential to resolve the V2A-T2A competition and foster a cooperative dynamic.
To further emphasize that this bridging effect is contingent on data quality, we provide a comparison in row 4. Here, we use standard-quality data (TA/VTA), where captions were generated by Gemini-2.5 using only the audio modality. Although the VT2A task is present, the poor V-T-A alignment fails to resolve the competition, and performance is still severely compromised (e.g., T2A FAD 1.13), far underperforming the SoundAtlas-driven model. This comparison proves that it is not merely the VT2A task, but the high-fidelity alignment of the bridge data, that is essential. This high quality enables data efficiency: the T2A ratio can be dropped to $\pi_{T2A}=0.1$ while achieving SOTA T2A performance, mitigating resource contention as designed.
Necessity of the Progressive Three-Stage Schedule.
Next, we demonstrate the necessity of our full progressive schedule in Table 6. We compare our full S1 $\rightarrow$ S2 $\rightarrow$ S3 pipeline against three baselines, all trained for the same total steps on SoundAtlas data. First, comparing the S2 only model with the S1 $\rightarrow$ S2 model confirms the value of the Stage 1 generative prior. Without S1, the S2 only model fails to converge well, showing poor quality (T2A FAD 1.22, V2A FAD 0.68). The S1 $\rightarrow$ S2 model, benefiting from the pretraining, significantly boosts generation quality (T2A FAD 0.94, V2A FAD 0.57) and resolves the Cross-Task Competition. However, this model still suffers from Intra-Task Competition (modality bias), as evidenced by its weaker A-V synchronization (V2A DS 0.49). Second, we validate our crucial hypothesis that Stage 3 must be decoupled. The S1 $\rightarrow$ [S2+S3] baseline, which merges the S3 robustness augmentations directly into S2, destabilizes the fragile optimization process. While it maintains A-V synchronization (V2A DS 0.47), introducing these augmentations prematurely harms the generative quality achieved in S2, leading to a clear degradation in FAD/FD scores (e.g., V2A FAD 0.60, VT2A FAD 0.61).
[…]
Method FAD KL LA-CLAP Win Rate
S1 $\rightarrow$ S2 0.97 1.46 0.31 46.8%
S1 $\rightarrow$ S2 $\rightarrow$ S3 0.85 1.39 0.32 53.2%
7 Conclusion
In this work, we addressed the foundational challenges of unified video-text-to-audio (VT2A) generation: data scarcity and inter-task competition. We introduce a three-part contribution: SoundAtlas, the first large-scale, human-expert-level audio caption dataset; Omni2Sound, a unified model featuring a three-stage progressive schedule to resolve task competition; and VGGSound-Omni, a comprehensive benchmark for unified VT2A evaluation. Our experiments demonstrate that this approach effectively resolves inter-task and intra-task competition and enables Omni2Sound to achieve unified state-of-the-art performance.
[…]
Supplementary Material
Overview
This document provides technical details, evaluation protocols, and extended experimental analyses. We begin with the Cost Analysis in Section A, validating SoundAtlas as a scalable and cost-effective pipeline. We then provide the exact Audio Caption Prompt Instructions in Section B, followed by detailed Evaluation Protocols to compare the quality of Audio Caption Datasets in Section C and the detailed construction of the Off-Screen Benchmark Track in Section D. Furthermore, we demonstrate the model’s Generalization Capabilities on third-party benchmarks in Section E and elaborate on the User Study in Section F. Section G outlines the Implementation Details, including model configurations and training data composition. Section H defines the Objective Evaluation Metrics on Generation Audio used throughout the paper. Qualitative results can be found in static HTML file.
Appendix A Cost Analysis on Audio Captioning
While Gemini 2.5 Pro [team2023gemini] represents a milestone as a native multimodal foundation model, utilizing it directly for large-scale video-grounded audio captioning proves economically unsustainable. As quantified in Table 6, using Gemini’s standard API pricing, a naive implementation—processing raw video frames alongside audio (V+AV+A)—incurs a prohibitive expenditure of \$10,275 USD per 1M samples. This figure is derived from the token consumption of a 10-second sample: the input aggregates to 3,820 tokens (comprising 1,000 instruction, 320 audio, and 2,500 visual tokens), while the full chain-of-thought generation requires $\sim$550 output tokens. Crucially, this naive approach suffers from an inherent visual bias, as shown in Figure 1 in main paper.
To address these challenges, our SoundAtlas pipeline employs three strategic optimizations. First, we implement Vision-to-Language Compression. This strategy replaces expensive raw video with a concise video caption $c_{v}$, eliminating the large $\sim$2,500 token visual overhead (Table 6, Row 2) and effectively mitigating the visual modality bias. Second, we enforce Restricted Reasoning, capping the generation output at $\sim$160 tokens (Table 6, Row 3). Finally, we utilize a Junior-Senior Agent Handoff that defaults to the cost-effective Flash model $G_{\text{junior}}$ for the majority of samples, reserving the Senior agent ($G_{\text{senior}}$) solely for complex cases. As shown in Table 6, while the standalone Flash model offers the lowest theoretical cost (\$1,026), our hybrid pipeline strikes a balance between quality and efficiency, reducing the initial expenditure of \$10,275 to approximately \$2,000 per million samples.
Table 6: Cost Analysis on Audio Captioning with Gemini 2.5. We compare the inference costs for processing one million 10-second samples. The table demonstrates a step-by-step ablation path: removing raw video (Row 2), restricting reasoning with vision-to-language compression (Row 3), and switching to the Flash model (Row 4) progressively reduces costs from \$10,275 to \$1,026.
Model Configuration Input Modality Input Token Num. Output Token Num. Est. Cost (USD / 1M Samples)
Gemini 2.5 Pro (Thinking-Full) T + V + A 3,820 550 \$10,275.00
Gemini 2.5 Pro (Thinking-Full) T + A 1,340 550 \$7,175.00
Gemini 2.5 Pro (Thinking-128) T + A 1,340 160 \$3,275.00
Gemini 2.5 Flash (Thinking-128) T + A 1,340 160 \$1,026.00
Table 7: Comparison of the generation performance on unified VT2A models and T2A models on Audiocaps test set.
Method KL FD FAD PQ LA-CLAP
AudioLDM 2-L [liu2024audioldm2] 1.73 34.21 2.26 5.93 0.24
TANGO 2 [majumder2024tango2] 1.19 15.92 3.17 5.82 0.35
Make-An-Audio 2 [huang2023makeanaudio2_placeholder] 1.38 15.34 1.46 5.64 0.25
GenAU-Large [haji-ali2024genau] 1.42 16.92 1.32 5.52 0.26
MMAudio [cheng2025mmaudio] 1.43 13.78 2.92 5.30 0.29
AudioX [tian2025audiox] 1.55 17.10 2.65 5.81 0.31
[…]
Appendix B Audio Caption Prompt Instructions
As illustrated in Figure 5, we present the audio captioning system prompt employed in our agentic annotation pipeline to construct the SoundAtlas dataset.
Appendix C Audio Caption Dataset Comparison
We provide the detailed scoring process for both MLLM-as-a-judge and Human Expert Evaluation on different audio caption datasets in Table 1 and 2 of main paper. The evaluation methodology consists of two stages: (1) absolute scoring based on the specific linguistic criteria defined below, and (2) a comparative win-rate calculation derived from these scores.
Subjective Evaluation Protocol.
We formulate a standardized scoring protocol for both MLLM and human evaluators, focusing on two distinct dimensions of modality alignment.
1. Semantic Alignment (MOS-S, Scale 1-4). This metric assesses both Accuracy (factuality of sound events) and Detail (precision of adjectives). The scale is defined as: (1) Factually incorrect/Brief; (2) Mostly incorrect/Brief; (3) Minor errors/Detailed (but visually redundant); and (4) Error-free and Detailed (strictly audio-centric).
2. Temporal Alignment (MOS-T, Scale 1-3). This evaluates whether the chronological order of described events matches the audio stream. The scale ranges from (1) Disordered, (2) Partially Correct, to (3) Perfectly Ordered. Samples with constant or stationary sounds (lacking distinct temporal events) are marked as N/A and excluded from this metric.
Human Evaluation Setup. To complement and validate our automated evaluation, we conducted a dedicated human expert evaluation based on the aforementioned protocol. We randomly sampled a subset of 100 instances from the evaluation corpus used in the MLLM-as-a-judge benchmark. We recruited five expert annotators with professional backgrounds in audio-visual analysis to assess these samples independently. To ensure robustness and mitigate individual bias, the final score for each item is derived by calculating the average rating across the five evaluators. For reference, the user study interface is illustrated in Figure 7.
Win Rate Calculation. We adopt a general pairwise comparison paradigm. For each evaluation set, a target model is compared against an opposing method. The Mean Win Rate (MWR) for any given model is derived by aggregating the outcomes of all its pairwise comparisons:
$\text{MWR}=\frac{N_{\text{win}}+0.5\times N_{\text{tie}}}{N_{\text{total}}}$ (1)
where $N_{\text{win}}$, $N_{\text{tie}}$, and $N_{\text{total}}$ denote the number of wins (scoring 1.0), ties (scoring 0.5), and total pairwise comparisons involving that model, respectively.
Appendix D Off-Screen Track of VGGSound-Omni
We introduce a dedicated Off-Screen Audio-Generation Track of VGGSound-Omni. This subset specifically evaluates the model’s capacity to handle non-depicted audio sources and is constructed through two distinct pipelines: (i) a Natural Off-screen Events subset sourced from the original test set; and (ii) a Synthetic Music subset focusing on background music (BGM) generation.
Natural Off-screen Events.
We construct the Natural Events subset by identifying VGGSound clips that inherently contain off-screen audio cues. The curation involves a rigorous three-step filtering pipeline. First, regarding Metadata & Modality, we ensure acoustic purity by excluding samples with pre-existing background music, static imagery, or voice-overs. Crucially, we filter out videos containing vision-only (“V”) labels, retaining only those with Audio-Visual (“AV”) or Audio-only (“A”) modalities. Second, for Complexity & Consistency, we limit scene complexity to a maximum of 6 labels. To capture "natural" off-screen scenarios, we filter based on the AV Ratio—defined as the proportion of “AV” labels relative to the total label count. We explicitly select samples where this ratio falls within $[0.25,0.80]$, ensuring that the audio content is not perfectly aligned with the visual stream (i.e., low A-V correspondence). Finally, we apply Distribution Balancing to mitigate the over-representation of common classes, restricting the proportion of speech to 20%.
(a) Subjective Evaluation Results on VT2A Task
(b) Subjective Evaluation Results on V2A Task
Figure 4: Subjective Evaluation Results on VGGSound-Omni. We report Mean Opinion Scores (MOS) on a 1-5 scale across three dimensions: Acoustic Quality (MOS-Q), Semantic Alignment (MOS-S), and Temporal Alignment (MOS-T). Omni2Sound consistently outperforms competitive baselines (AudioX, MMAudio, HunyuanVideo-Foley, Frienren-V2A) across all perceptual metrics on both VT2A and VT2A tasks, validating its superior generation fidelity and alignment.
Synthetic Music Augmentation.
To address the high demand for Background Music (BGM) generation, we create a Synthetic Music subset by mixing semantically aligned MusicCaps [agostinelli2023musiccaps] clips into a pool of high-fidelity videos. This process follows a two-stage procedure. In the Base Selection stage, we first select a "clean" video pool by strictly requiring a 100% AV label ratio and filtering for high alignment (ImageBind $\geq 0.30$, Desync <0.55<0.55), ensuring all original acoustic events are visually manifest. Subsequently, during Semantic Mixing, we augment these videos with background music tracks. To guarantee semantic coherence, we utilize GPT to retrieve the most congruent music track from a random candidate batch of 50 samples based on the video context. Ground-truth captions are updated to reflect this acoustic addition.
Comparison with Concurrent Work.
We acknowledge the pioneering work of VinTAGe-Bench [bain2025vintage] in synthetic robustness evaluation. However, the off-screen subset of our VGGSound-Omni benchmark extends this direction in three critical dimensions. First, in terms of Realism, by leveraging VGGSounder [zverev2025vggsounder] metadata, our natural subset is primarily sourced from real-world off-screen audio events rather than relying solely on synthetic mixes. Second, regarding Scale, our benchmark is significantly larger, providing 1,613 evaluation items compared to the 212 basic videos of VinTAGe-Bench. Third, regarding Scope, we include a dedicated Synthetic Music (BGM) track, addressing a critical, high-demand scenario often overlooked in standard environmental sound benchmarks.
Appendix E Generalization on Third-Party Benchmarks.
To further validate our model’s generalization and mitigate potential biases from our self-constructed benchmark, we evaluate it on the Kling-Audio-Eval [wang2025audiogenomni] and Audiocaps test set [kim2019audiocaps]. In Table 4, on the Kling-Audio-Eval benchmark, Omni2Sound remains highly competitive, despite a significant data scale and distribution gap (our YouTube-sourced SoundAtlas vs. Kling’s professional video/Foley). While HunyuanVideo-Foley [chen2024hunyuanfoley] leads on several metrics, this is expected given its massive 100k-hour internal dataset, which is tens of times larger than our SoundAtlas filter derived from VGGSound and AudioSet. Nevertheless, Omni2Sound consistently outperforms all other strong baselines (e.g., MMAudio, AudioX, and ThinkSound) across V2A and VT2A tasks, demonstrating strong generalization as the SOTA or second-best method. In Table 7, on the Audiocaps test set, we compare Omni2Sound against specialized SOTA T2A models. The results show our unified model achieves top-tier performance, attaining the best scores in key distribution metrics ($\mathrm{KL}$, $\mathrm{FD}$) and semantic alignment ($\mathrm{CLAP}=0.36$), while remaining highly competitive in audio quality (PQ) and the $\mathrm{FAD}$ metric.
Appendix F User Study
We conduct a comprehensive user study on the VGGSound-Omni benchmark to validate Omni2Sound against top baselines (four methods in total). Given the density of comparisons involved, we structure VT2A and V2A as independent evaluation tracks to mitigate evaluator fatigue. We recruit a total of 16 expert evaluators, who are evenly distributed across the two independent tasks. Each participant evaluates 20 random samples (80 comparisons) within their assigned track. Samples from the same source are grouped with randomized method order to maintain blinding. In total, 1280 responses per metric are collected.
Subjective Evaluation Metrics.
Our final evaluation utilizes a multi-dimensional Mean Opinion Score (MOS) protocol, where expert human evaluators assess the generated audio across three distinct criteria. All scores are normalized to a 5-point Likert scale (1: Poor/Misaligned; 5: Excellent/Perfectly Aligned).
  • MOS-Q: Acoustic Fidelity (Quality). This metric assesses the intrinsic acoustic quality and perceptual realism of the generated sound, independent of the conditioning inputs. Evaluators focus on auditory naturalness, clarity, and the absence of technical artifacts (e.g., distortion, noise, mixing comfort).
  • MOS-S: Semantic Consistency (Alignment). This quantifies the perceptual fidelity between the content of the generated audio and the semantic information conveyed by the conditioning modalities (video frames and textual captions). Evaluation centers on whether the generated sound event’s category and characteristics logically correspond to the depicted visual and textual context.
  • MOS-T: Temporal Synchronization (Alignment). This assesses the temporal accuracy of the acoustic events against the visual stream. Evaluators specifically check the precision of sound onset, offset, and duration, ensuring tight synchronization with the corresponding visual event timing.
The results, summarized in Figure 4, demonstrate that Omni2Sound outperforms all baselines across the three subjective metrics: MOS-Q, MOS-S, and MOS-T on both VT2A and V2A tasks. This strong alignment between human preference in Figure 4 and the objective metrics presented in Table 3 in main paper validates the effectiveness of our proposed data construction and training pipeline. For reference, the user study interface is illustrated in Figure 7.
Appendix G Implementation Details
Model Configuration.
Following Stable Audio [evans2024stable], our diffusion model adopts a Diffusion Transformer (DiT) architecture within a Latent Diffusion Model (LDM) paradigm. The diffusion backbone consists of a DiT with 24 layers, 24 attention heads, and a hidden dimension of 1536. We employ cross-attention mechanisms to inject semantic conditions (e.g., FLAN-T5 and CLIP embeddings) and Adaptive Layer Normalization (AdaLN) to integrate temporal signals, as detailed in Section 4.1. Both the conditional token dimension and the global condition embedding dimension are 1024. Finally, for audio compression, we train a Variational Autoencoder (VAE) from scratch based on the wav Audio VAE architecture [evans2024stable], operating at a 16kHz sampling rate. With strides of $[4,4,4,10]$, the encoder achieves a total downsampling ratio of 640, mapping mono waveforms into a compact 64-dimensional latent space. To ensure high-fidelity reconstruction, we utilize Snake activations throughout the network.
Training Data.
For T2A backbone pre-training, we use a large-scale corpus comprising the train set of audio datasets such as AudioCaps [kim2019audiocaps], WavCaps [mei2023wavcaps], Clotho [drossos2020clotho], AudioSet [gemmeke2017audio], VGGSound [chen2020vggsound], FSD50k [fonseca2022fsd50k], as well as music datasets including MSD [bertin2011million] and FMA [defferrard2016fma]. All audio signals are standardized to a mono-channel format at 16kHz. To accommodate fixed-size diffusion inputs, we normalize clips to a uniform 10-second duration: samples exceeding this length undergo right cropping, while shorter samples are right-padded with silence.
Subsequently, the model is fine-tuned for unified multimodal tasks using our proposed SoundAtlas. Constructed following the pipeline detailed in Section 5, this dataset comprises 470k high-quality V-A-T pairs, sourced from 140k VGGSound and 330k AudioSet samples. Notably, the AudioSet subset is strictly curated: starting from the original 2M corpus, we first applied a preliminary filtration to exclude all speech- and music-related categories, resulting in a candidate pool of 450k sound samples. These candidates then underwent our A-V consistency routing and verification pipeline to yield the final 330k high-fidelity pairs. For T2A task fine-tuning, we augment the training with T-A pairs from SoundAtlas as well as a high-fidelity subset of the pre-training corpus, filtered by strict quality thresholds: requiring a CLAP score greater than 0.35 and a PQ score exceeding 6.0.
Appendix H Objective Evaluation Metrics.
We implement our objective evaluation metrics using the standardized AV-benchmark toolkit [hkchengrex2024avbenchmark]. All samples are generated under the same video and text conditions and evaluated in 8-second clips, following previous work [cheng2025mmaudio]. Following common practice [liu2023audioldm], we assess the quality of the generation in four critical dimensions. For Distribution Matching, we measure the similarity in feature distribution between generated and ground-truth audio. We compute the Fréchet Distance using the VGGish (FAD) [hershey2017cnn] and PaSST ($\mathrm{FD}_{\mathrm{PaSST}}$) [koutini2022passt] embeddings, as well as the Fréchet Audio Distance using PANNs (FD) [kong2020panns]. We also report the Kullback-Leibler divergence using PANNs (KL) and PaSST ($\mathrm{KL}_{\mathrm{PaSST}}$) classifiers. For Audio Quality, we assess the quality of the generation using the Inception Score [salimans2016improved], calculated with both the PANNs (IS) and PaSST ($\mathrm{IS}_{\mathrm{PaSST}}$) classifiers. For Semantic Alignment, we evaluate text-audio consistency using LAION CLAP (CLAP) [elizalde2023clap] and Microsoft CLAP (MS-CLAP) [wu2023large] scores, and video-audio alignment using ImageBind score (IB) [girdhar2023imagebind] as cosine similarity between video and audio embeddings. Finally, for Temporal Alignment, we assess audio-visual synchrony using the DS metric predicted by Synchformer [iashin2024synchformer].
Audio Captioning Instruction for SoundAtlas Roles and Tasks You are an experienced audio content analyst skilled in describing soundscapes through detailed, multi-dimensional natural language. Given an audio clip (a) and its corresponding video descriptions ($T_{v}$), identify and describe all relevant auditory elements in chronological order, then write a rich audio description that faithfully and dynamically reflects the scene. Annotation Dimensions 1. Primary Sound Information Humans/Animals: speech (talking, shouting), movements (footsteps). Note: Do not transcribe words/lyrics; describe voice characteristics. Objects: traffic, office sounds, battlefield, tools. Characteristics: Gender/age, language, quantity (monologue/turn-taking), emotional tone, voice qualities. 2. Background Sounds (if present) Natural (wind, rain) or Artificial (city noise, crowds). Briefly specify the environment if necessary. 3. Music (if present) Style/genre, rhythmic features, emotional tone, atmosphere. Identifiable instruments and effects (harmonies, reverb). 4. Detailed Descriptors Changes in volume/speed/intensity. Narrative functions. Detailed duration, spatial distance, pitch, timbre, texture. Important Guidelines 1. Avoid Redundancy: Identify sources once unless they change significantly. Keep it concise. 2. Prioritize the Audio: Use video description only to clarify ambiguous sounds. If a sound isn’t audible, don’t describe it. 3. Avoid Hallucinated Sounds: Only describe perceptible sounds. Avoid describing artifacts (e.g., "high-pitched squeal" from edits). Output Format Integrate elements into one or few sentences following these rules: Language: English. Structure: No lists or bullet points. Length: Max 40 words. Concise but detailed. Temporal Order: Chronological (e.g., "first", "then", "suddenly"). Style: Natural, objective, context-sensitive. Focus on what is heard.   Examples Example 1 (General): Input: [High-pitched mechanical whirring with periodic thuds]
Video Caption: "Laundromat with washing machines and dryers running"
Output: Washing machines whir at high speed while dryers tumble clothes with periodic rhythmic thuds. Water drains intermittently as cycles complete and doors slam shut.
Example 2 (Anti-hallucination): Input: [Guitar strumming and melody]
Video Caption: "Musician performing with piano and guitar on stage"
Output: Acoustic guitar plays melodic fingerpicking patterns with clear, resonant tones. (Piano is omitted as it is not audible).
Figure 5: Audio Captioning Instruction for SoundAtlas.
Figure 6: User study interface for human evaluation across different audio generation models.
Figure 7: User study interface for human evaluation across different automatic audio captioning datasets.

### v2

@@
 Early audio generation models typically rely on unimodal conditioning.
-Text-to-Audio (T2A) [kreuk2022audiogen, liu2023audioldm, evans2024stableaudioopen, ghosal2023instruction] offers strong semantic fidelity and generalization but lacks dense temporal control.
-Conversely, Video-to-Audio (V2A) [luo2023diff-foley, wang2024frieren, zhang2024foleycraft, chen2024videoguided] ensures fine-grained temporal synchronization with video, yet suffers from weak reasoning in complex scenes and unfaithful generation (e.g., unexpected music or speech) [bain2025vintage, yu2024reasonaudio].
-To address this, recent Video-Text-to-Audio (VT2A) methods [chen2024hunyuanfoley, liu2025thinksound, cheng2025mmaudio, tian2025audiox, bain2025vintage] jointly condition on video and text.
-While VT2A achieves strong both semantic understanding and temporal alignment, its reliance on simultaneous inputs constrains its applicability [tian2025audiox].
-Crucially, most VT2A systems lack robustness [liu2025thinksound, chen2024hunyuanfoley, bain2025vintage], degrading sharply under missing-modality conditions (video-only or text-only).
+Text-to-Audio (T2A) [27, 30, 13, 16] offers strong semantic fidelity and generalization but lacks dense temporal control.
+Conversely, Video-to-Audio (V2A) [33, 47, 53, 6] ensures fine-grained temporal synchronization with video, yet suffers from weak reasoning in complex scenes and unfaithful generation (e.g., unexpected music or speech) [29, 21].
+To address this, recent Video-Text-to-Audio (VT2A) methods [40, 32, 7, 43, 29] jointly condition on video and text.
+While VT2A achieves strong both semantic understanding and temporal alignment, its reliance on simultaneous inputs constrains its applicability [43].
+Crucially, most VT2A systems lack robustness [32, 40, 29], degrading sharply under missing-modality conditions (video-only or text-only).
 These constraints motivate a unified framework natively supporting VT2A, V2A, and T2A within a single model.
@@
 Recent work has begun to advance this unified approach.
-MMAudio [cheng2025mmaudio] introduces a multimodal joint training framework to improve V2A generation, optionally conditioning on text using large-scale text–audio pairs.
-Moreover, AudioX [tian2025audiox] enhanced flexibility by supporting broader modality combinations.
+MMAudio [7] introduces a multimodal joint training framework to improve V2A generation, optionally conditioning on text using large-scale text–audio pairs.
+Moreover, AudioX [43] enhanced flexibility by supporting broader modality combinations.
 Despite this progress, two challenges in the unified VT2A framework remain underexplored.
@@
 First, there is a scarcity of high-quality audio captions that are well-aligned with both audio and video cues.
-Most unified or specialized VT2A studies create their (V, T, A) training triplets by pairing videos (V) and their audio (A) with captions (T) generated solely from the audio [tian2025audiox, chen2024hunyuanfoley].
+Most unified or specialized VT2A studies create their (V, T, A) training triplets by pairing videos (V) and their audio (A) with captions (T) generated solely from the audio [43, 40].
 However, this approach introduces severe semantic conflict in the multimodal training data (see Figure 1): a frequent mismatch between the visual content and the (audio-only) text caption.
@@
 an electric drill).
-This fundamental ambiguity is then exacerbated by the limited capabilities of earlier audio-language models, which are prone to severe hallucinations (e.g., omissions and mislabels) [chen2025detecting].
+This fundamental ambiguity is then exacerbated by the limited capabilities of earlier audio-language models, which are prone to severe hallucinations (e.g., omissions and mislabels) [5].
 In our preliminary experiments, we found these modality conflicts caused by mismatches between V-T conditions directly lead to unstable convergence and a significant degradation in audio faithfulness.
@@
 (1) Cross-Task Competition.
-Prior work, notably MMAudio [cheng2025mmaudio], established that incorporating T-A pairs enhances the generalization and quality of V2A generation.
+Prior work, notably MMAudio [7], established that incorporating T-A pairs enhances the generalization and quality of V2A generation.
 However, training a unified model to excel at both V2A and T2A presents a significant challenge: as shown in our preliminary experiment (Table 5), this joint training introduces a severe T2A-V2A adverse trade-off, rooted in the heterogeneity between text and video modalities.
@@
 To address data scarcity, we first introduce SoundAtlas in Section 3, a large-scale, agent-generated multimodal audio-caption dataset.
-It augments the two largest audio datasets, VGGSound [chen2020vggsound] and AudioSet [gemmeke2017audio], providing semantically rich and temporally detailed captions that even surpass human-expert quality (Table 2).
-Built on current advanced multimodal foundation models (Gemini-2.5 Pro [team2023gemini] and Qwen-2.5-VL [qwen2025vl]), we develop a multi-turn, agentic annotation pipeline featuring a junior–senior agent handoff, vision-to-language compression, and post-hoc hallucination filtering.
+It augments the two largest audio datasets, VGGSound [4] and AudioSet [15], providing semantically rich and temporally detailed captions that even surpass human-expert quality (Table 2).
+Built on current advanced multimodal foundation models (Gemini-2.5 Pro [42] and Qwen-2.5-VL [28]), we develop a multi-turn, agentic annotation pipeline featuring a junior–senior agent handoff, vision-to-language compression, and post-hoc hallucination filtering.
 This pipeline delivers cost-controlled annotations while maintaining tight visual–audio–text (V–A–T) alignment and a markedly higher text-audio faithfulness than prior datasets.
-Interestingly, we find its quality is high enough to even correct human annotation errors in VGGSounder [zverev2025vggsounder].
+Interestingly, we find its quality is high enough to even correct human annotation errors in VGGSounder [54].
 Building on this dataset, we propose Omni2Sound in Section 4, a diffusion-based unified model supporting flexible input modalities while maintaining both audio-visual synchronization and high-fidelity generation.
@@
 It provides high-quality, human-verified annotations for all three tasks and introduces a challenging off-screen audio generation track.
-As a result, with a vanilla DiT [Peebles2023DiT] backbone, Omni2Sound achieves unified state-of-the-art performance across all three tasks against both unified and specialized models, showing high-fidelity audio quality, tight audio-visual synchronization, and excellent generation faithfulness.
+As a result, with a vanilla DiT [37] backbone, Omni2Sound achieves unified state-of-the-art performance across all three tasks against both unified and specialized models, showing high-fidelity audio quality, tight audio-visual synchronization, and excellent generation faithfulness.
 2 Related Works
 Audio Caption Dataset.
-Human-annotated benchmarks like AudioCaps [kim2019audiocaps] (46k) and Clotho [drossos2020clotho] (5k) offer high-quality alignment, but their limited scale, high cost, and lack of detail make them unsuitable for training modern, large-scale models.
+Human-annotated benchmarks like AudioCaps [24] (46k) and Clotho [10] (5k) offer high-quality alignment, but their limited scale, high cost, and lack of detail make them unsuitable for training modern, large-scale models.
 Automated pipelines emerged to address data scarcity.
-WavCaps [mei2023wavcaps] used LLMs to refine noisy web metadata (400k captions), and AudioSetCaps [lipping2024audiocaps] used ALMs+LLMs to extract and aggregate details from audio, speech, and music, significantly increasing data volume.
+WavCaps [36] used LLMs to refine noisy web metadata (400k captions), and AudioSetCaps [2] used ALMs+LLMs to extract and aggregate details from audio, speech, and music, significantly increasing data volume.
 As detailed in our Introduction, these audio-only methods suffer from high hallucination rates and can lead to cross-modal conflicts that destabilize VT2A training, stemming from the audio modality’s inherent ambiguity.
-Visually-enhanced (VE) annotation pipelines like Auto-ACD [han2024autoacd] and Sound-VECaps [yu2024soundvecaps] emerged to leverage visual cues for cross-modal constraint.
+Visually-enhanced (VE) annotation pipelines like Auto-ACD [41] and Sound-VECaps [51] emerged to leverage visual cues for cross-modal constraint.
 While promising, existing implementations adopt a separate-then-fuse pipeline: unimodal models extract separate textual cues (e.g., image captions, audio tags), which are then merged by a final LLM.
 This pipeline is suboptimal, as the LLM fuses lossy textual representations, not raw modalities, leading to the accumulation and amplification of unimodal hallucinations.
-While using native end-to-end multimodal models (e.g., Gemini [team2023gemini]) seems a natural solution, it also proves suboptimal.
+While using native end-to-end multimodal models (e.g., Gemini [42]) seems a natural solution, it also proves suboptimal.
 As we demonstrate in Section 3, this method faces prohibitive costs and a pervasive visual bias that prevents truly audio-centric captioning.
@@
 Unified Audio Generation Model.
-The audio generation paradigm is shifting towards unified, omni-modal frameworks, a trajectory initiated by MMAudio [cheng2025mmaudio].
+The audio generation paradigm is shifting towards unified, omni-modal frameworks, a trajectory initiated by MMAudio [7].
 While it integrated V2A and T2A, its approach was fundamentally V2A-centric, using T-A pairs merely as augmentation for V2A rather than optimizing T2A as a co-equal task.
-Subsequent works like AudioX [tian2025audiox] and AudioGen-Omni [wang2025audiogenomni] expanded this scope to more flexible modality combinations.
+Subsequent works like AudioX [43] and AudioGen-Omni [46] expanded this scope to more flexible modality combinations.
 However, these efforts often relied on brute-force data scaling (e.g., AudioX with over 9 million samples), which revealed inefficiencies and failed to yield proportional SOTA performance.
-Critically, these early models [cheng2025mmaudio, tian2025audiox, wang2025audiogenomni] largely overlooked the inherent cross-task competition stemming from co-training these diverse sub-tasks.
-UniFlow-Audio [xu2025uniflow] is the first to systematically address this by categorizing tasks into Time-Aligned (TA) and Non-Time-Aligned (NTA) classes and analyzing their competitive dynamics.
+Critically, these early models [7, 43, 46] largely overlooked the inherent cross-task competition stemming from co-training these diverse sub-tasks.
+UniFlow-Audio [50] is the first to systematically address this by categorizing tasks into Time-Aligned (TA) and Non-Time-Aligned (NTA) classes and analyzing their competitive dynamics.
 However, its analysis remains coarse-grained, failing to investigate the granular competition within the TA category (i.e., V2A vs.
@@
 3 SoundAtlas: V-A-T Data Construction
-Existing automated audio caption datasets often suffer from severe visual-audio-text (V-A-T) misalignment with high hallucination rates due to the limitations of early ALMs [lipping2024audiocaps, han2024autoacd, yu2024soundvecaps].
-While recent native multimodal foundation models like Gemini 2.5 [team2023gemini, MMAR:2025, Qwen3Omni:2025] offer strong capabilities, we find that a naive implementation—processing raw video-audio pairs directly—is suboptimal for audio caption dataset construction.
+Existing automated audio caption datasets often suffer from severe visual-audio-text (V-A-T) misalignment with high hallucination rates due to the limitations of early ALMs [2, 41, 51].
+While recent native multimodal foundation models like Gemini 2.5 [42, 34, 49] offer strong capabilities, we find that a naive implementation—processing raw video-audio pairs directly—is suboptimal for audio caption dataset construction.
 Specifically, it incurs prohibitive costs (approx.
@@
 A-V Consistency Routing.
-We first apply A-V Consistency Routing on raw video from AudioSet [gemmeke2017audio] and VGGSound [chen2020vggsound].
+We first apply A-V Consistency Routing on raw video from AudioSet [15] and VGGSound [4].
 This step is based on the core finding that visual cues are reliable for high-consistency A-V clips but act as distractors in low-consistency clips as shown in Figure 2.
@@
 LA-CLAP $\uparrow$ MS-CLAP $\uparrow$ LA-CLAP $\uparrow$ MS-CLAP $\uparrow$
-AudioSetCaps [lipping2024audiocaps] 0.330 0.397 0.351 0.421
-Sound-VECaps [yu2024soundvecaps] 0.370 0.425 - -
-Auto-ACD [han2024autoacd] 0.396 0.437 0.409 0.457
+AudioSetCaps [2] 0.330 0.397 0.351 0.421
+Sound-VECaps [51] 0.370 0.425 - -
+Auto-ACD [41] 0.396 0.437 0.409 0.457
 SoundAtlas (Ours) 0.447 0.485 0.461 0.502
 Table 2: Caption quality comparison via MLLM-as-a-judge and human evaluation, reporting the Mean Win Rate for Semantic (MWR-S) and Temporal (MWR-T) alignment.
-Human-Expert refers to the human-annotated captions from AudioCaps [kim2019audiocaps].
+Human-Expert refers to the human-annotated captions from AudioCaps [24].
 Method MLLM Evaluation Human Evaluation
 MWR-S $\uparrow$ MWR-T $\uparrow$ MWR-S $\uparrow$ MWR-T $\uparrow$
-Auto-ACD [han2024autoacd] 0.39 0.41 0.31 0.26
-Human-Expert [kim2019audiocaps] 0.36 0.51 0.46 0.55
+Auto-ACD [41] 0.39 0.41 0.31 0.26
+Human-Expert [24] 0.36 0.51 0.46 0.55
 SoundAtlas (Ours) 0.75 0.58 0.71 0.69
@@
 Second, it robustly mitigates cross-modal hallucinations by filtering the visual bias, providing only low-bias semantic context (e.g., "A man and a woman are standing…") rather than a misleading raw visual stream.
-Therefore, for samples $V$ routed to the A-V Enhanced Path, we use Qwen-2.5-VL [qwen2025vl] to analyze the video $V$ (without its audio $A$) and generate the textual representation $c_{v}=\text{Qwen}(V)$.
+Therefore, for samples $V$ routed to the A-V Enhanced Path, we use Qwen-2.5-VL [28] to analyze the video $V$ (without its audio $A$) and generate the textual representation $c_{v}=\text{Qwen}(V)$.
 Conversely, samples on the Audio-Only Path are assigned a null context.
@@
 Let the output caption be $c_{a}=G_{\text{junior}}(A,c_{v})$.
-This caption $c_{a}$ is then flagged if it (i) meets our complexity criteria (text-based heuristics to identify complex audio scenes), (ii) contains high-frequency hallucination phrases, or (iii) fails our differentiated CLAP [elizalde2023clap] check, $\text{CLAP}(c_{a},A)<\tau_{clap}$, where $\tau_{clap}$ is $0.35$ for general audio and $0.15$ for music.
+This caption $c_{a}$ is then flagged if it (i) meets our complexity criteria (text-based heuristics to identify complex audio scenes), (ii) contains high-frequency hallucination phrases, or (iii) fails our differentiated CLAP [11] check, $\text{CLAP}(c_{a},A)<\tau_{clap}$, where $\tau_{clap}$ is $0.35$ for general audio and $0.15$ for music.
 Flagged tasks are escalated to the Senior agent, $G_{\text{senior}}$ (Gemini 2.5 Pro).
@@
 Finally, all generated captions $c_{a}$ undergo a two-stage verification.
-First, a CLAP (T-A) filtering model [elizalde2023clap] ensures high Text-Audio faithfulness; captions where $\text{CLAP}(c_{a},A)<\tau_{verify}$ are discarded.
+First, a CLAP (T-A) filtering model [11] ensures high Text-Audio faithfulness; captions where $\text{CLAP}(c_{a},A)<\tau_{verify}$ are discarded.
 Second, for captions from the A-V Enhanced Path ($c_{v}\neq\emptyset$), an A-V-T Verifier, $V_{\text{AVT}}$, ensures $c_{a}$ is a reasonable acoustic inference given $c_{v}$.
-Captions that pass all filters are accepted into the final dataset $\mathcal{D}_{\text{SoundAtlas (Ours)}}$, which augments VGGSound [chen2020vggsound] and AudioSet [gemmeke2017audio] datasets with human-expert-level audio captions.
+Captions that pass all filters are accepted into the final dataset $\mathcal{D}_{\text{SoundAtlas (Ours)}}$, which augments VGGSound [4] and AudioSet [15] datasets with human-expert-level audio captions.
 3.1 Comparison with Existing Pipeline
-We compare SoundAtlas against other automated pipelines [lipping2024audiocaps, han2024autoacd, yu2024soundvecaps] on high audio-visual consistency subsets sourced from AudioSet and VGGSound, where ImageBind score $s_{ib}>0.30$.
+We compare SoundAtlas against other automated pipelines [2, 41, 51] on high audio-visual consistency subsets sourced from AudioSet and VGGSound, where ImageBind score $s_{ib}>0.30$.
 As shown in Table 1, SoundAtlas significantly outperforms all competitors on both LA-CLAP and MS-CLAP scores, demonstrating superior text-audio alignment.
-Additionally, we conduct a fine-grained MLLM-as-a-judge (Gemini 3.0 Pro [team2023gemini]) evaluation on the intersection of AudioCaps and all compared datasets [kim2019audiocaps].
+Additionally, we conduct a fine-grained MLLM-as-a-judge (Gemini 3.0 Pro [42]) evaluation on the intersection of AudioCaps and all compared datasets [24].
 As shown in Table 2, SoundAtlas achieves a substantially higher mean win rate in semantic alignment (MWR-S) and temporal alignment (MWR-T) than both the strongest baseline (Auto-ACD) and the Human-Expert annotations, across both semantic and temporal alignment.
@@
 4.1 Foundation Model Architecture
-We adhere to a principle of simplicity and scalability, adopting a standard Diffusion Transformer (DiT) backbone [Peebles2023DiT] conditioned on latent features from a pre-trained audio VAE [evans2024fast].
+We adhere to a principle of simplicity and scalability, adopting a standard Diffusion Transformer (DiT) backbone [37] conditioned on latent features from a pre-trained audio VAE [12].
 As shown in Figure 3, the backbone is conditioned on multimodal inputs using a decoupled injection approach, which is separated into two distinct branches: (1) Semantic Branch (What) and (2) Temporal Branch (When).
-To capture global semantic context, we concatenate text embeddings from Flan-T5 [chung2024scaling] ($F_{t}$) and visual features from CLIP [radford2021learning] ($F_{v}$, sampled at 8 fps) along the temporal dimension, which are then injected via cross-attention layers.
+To capture global semantic context, we concatenate text embeddings from Flan-T5 [8] ($F_{t}$) and visual features from CLIP [38] ($F_{v}$, sampled at 8 fps) along the temporal dimension, which are then injected via cross-attention layers.
 Crucially, this design allows for flexible unimodal generation (V2A or T2A) by simply omitting the absent modality without requiring padding constraints.
-For the Temporal Branch, to ensure fine-grained synchronization, we follow [cheng2025mmaudio] to utilize a Synchformer [Viertola2024Synchformer] to extract dense visual-temporal features ($F_{s}$) and then inject it globally via Adaptive Layer Normalization (AdaLN).
-This decoupled architecture effectively (1) achieves the flexibility of multi-condition frameworks like AudioX [cheng2025mmaudio], supporting extensible conditions without architectural modification; and (2) maintains precise temporal alignment comparable to MMAudio [cheng2025mmaudio] (powered by its well-designed MM-DiT architecture).
+For the Temporal Branch, to ensure fine-grained synchronization, we follow [7] to utilize a Synchformer [23] to extract dense visual-temporal features ($F_{s}$) and then inject it globally via Adaptive Layer Normalization (AdaLN).
+This decoupled architecture effectively (1) achieves the flexibility of multi-condition frameworks like AudioX [7], supporting extensible conditions without architectural modification; and (2) maintains precise temporal alignment comparable to MMAudio [7] (powered by its well-designed MM-DiT architecture).
 4.2 Three-stage Progressive Multi-task Training
@@
 We first conduct standalone T2A pretraining on large-scale text-audio pairs without a quality filter.
-Following the latent diffusion framework [stableaudio2024, Peebles2023DiT], our DiT backbone (Section 4.1) is trained to progressively denoise a noisy latent $z_{t}$ at timestep $t$, conditioned on text embeddings $H_{c}$.
+Following the latent diffusion framework [13, 37], our DiT backbone (Section 4.1) is trained to progressively denoise a noisy latent $z_{t}$ at timestep $t$, conditioned on text embeddings $H_{c}$.
 The model, $\epsilon_{\theta}$, is optimized via the simple L2 loss:
@@
 Table 3: Comparison on VGGSound-Omni benchmark: Omni2Sound against SOTA models on T2A, V2A, and VT2A tasks.
-The w/ Video-LLaMA caps row evaluates Omni2Sound’s generalization to unseen captions generated by Video-LLaMA [zhang2023video-llama].
+The w/ Video-LLaMA caps row evaluates Omni2Sound’s generalization to unseen captions generated by Video-LLaMA [52].
 Task Method Distribution Matching Audio Quality Modality Alignment
 KL FD FAD $\mathrm{\textbf{FD}_{\textbf{PaSST}}}$ PQ IS DS IB MS-CLAP
-T2A AudioX [tian2025audiox] 1.68 9.04 1.42 109.94 6.37 15.15 - - 0.49
-MMAudio [cheng2025mmaudio] 1.92 8.62 1.63 101.66 5.84 14.30 - - 0.50
+T2A AudioX [43] 1.68 9.04 1.42 109.94 6.37 15.15 - - 0.49
+MMAudio [7] 1.92 8.62 1.63 101.66 5.84 14.30 - - 0.50
 […]
     w/ Video-LLaMA caps 1.60 6.92 1.23 83.91 6.38 16.01 - - 0.51
-V2A V-AURA [viertola2024vaura] 2.28 16.43 2.34 245.25 5.74 10.82 0.69 0.28 0.32
-Frieren [wang2024frieren] 2.73 12.13 1.23 123.75 5.82 11.32 0.86 0.21 0.31
-AudioX [tian2025audiox] 2.96 12.73 1.42 121.82 6.17 13.34 1.22 0.24 0.34
-MMAudio [cheng2025mmaudio] 2.11 5.65 0.81 69.33 5.72 11.85 0.48 0.28 0.43
+V2A V-AURA [45] 2.28 16.43 2.34 245.25 5.74 10.82 0.69 0.28 0.32
+Frieren [47] 2.73 12.13 1.23 123.75 5.82 11.32 0.86 0.21 0.31
+AudioX [43] 2.96 12.73 1.42 121.82 6.17 13.34 1.22 0.24 0.34
+MMAudio [7] 2.11 5.65 0.81 69.33 5.72 11.85 0.48 0.28 0.43
 […]
-VT2A ThinkSound ($w/o$. CoT) [liu2025thinksound] 1.60 7.41 1.10 116.08 6.21 11.73 0.53 0.26 0.43
-HunyuanVideo-Foley [chen2024hunyuanfoley] 1.74 10.02 2.36 100.53 6.18 11.58 0.57 0.32 0.45
-AudioX [tian2025audiox] 1.59 8.29 1.24 103.37 6.17 14.94 1.23 0.26 0.49
-MMAudio [cheng2025mmaudio] 1.63 5.28 0.91 68.44 5.84 13.44 0.49 0.29 0.49
+VT2A ThinkSound ($w/o$. CoT) [32] 1.60 7.41 1.10 116.08 6.21 11.73 0.53 0.26 0.43
+HunyuanVideo-Foley [40] 1.74 10.02 2.36 100.53 6.18 11.58 0.57 0.32 0.45
+AudioX [43] 1.59 8.29 1.24 103.37 6.17 14.94 1.23 0.26 0.49
+MMAudio [7] 1.63 5.28 0.91 68.44 5.84 13.44 0.49 0.29 0.49
 […]
@@
 A significant challenge in evaluating unified Video-Text-to-Audio (VT2A) models is the absence of a comprehensive benchmark.
-The VGGSound test set [chen2020vggsound] only provides sparse event labels and lacks detailed captions.
-Although recent work like VGGSounder [zverev2025vggsounder] significantly improved this by correcting and adding crucial modality labels (e.g., A, V, AV) for fidelity evaluation, it still lacks human-expert-level captions.
+The VGGSound test set [4] only provides sparse event labels and lacks detailed captions.
+Although recent work like VGGSounder [54] significantly improved this by correcting and adding crucial modality labels (e.g., A, V, AV) for fidelity evaluation, it still lacks human-expert-level captions.
 To address this gap, we construct VGGSound-Omni, a new multi-track benchmark derived from the original VGGSound test set, designed for both standard unified and specialized off-screen VT2A tasks evaluations.
@@
 We first generated an initial caption using our agentic pipeline (Section 3).
-We then systematically validated this output via an AI-assisted verification workflow: GPT-5 [openai2025gpt5intro] was tasked to act as an auditor, checking if our captions semantically covered all the “A” and “AV” labels from VGGSounder [zverev2025vggsounder].
+We then systematically validated this output via an AI-assisted verification workflow: GPT-5 [openai2025gpt5intro] was tasked to act as an auditor, checking if our captions semantically covered all the “A” and “AV” labels from VGGSounder [54].
 Samples flagged with a mismatch were routed for targeted human verification.
@@
 Complementing the primary set, we construct a challenging off-screen track (1,048 items).
-We curated this subset from two sources: (i) Natural events, filtered from VGGSound for low A-V correspondence (via IB-Score [girdhar2023imagebind] and Desync-Score [cheng2025mmaudio]) while excluding background speech; and (ii) Synthetic music, formed by mixing aligned background clips from MusicCaps [agostinelli2023musiccaps].
+We curated this subset from two sources: (i) Natural events, filtered from VGGSound for low A-V correspondence (via IB-Score [17] and Desync-Score [7]) while excluding background speech; and (ii) Synthetic music, formed by mixing aligned background clips from MusicCaps [1].
 More Details are provided in Appendix D.
@@
 KL FD FAD $\mathrm{\textbf{FD}_{\textbf{PaSST}}}$ PQ IS DS IB LA-CLAP
-T2A AudioX [tian2025audiox] 2.73 19.43 3.32 171.60 5.98 12.15 - - 0.28
-MMAudio [cheng2025mmaudio] 2.54 11.25 5.07 142.71 5.54 9.28 - - 0.28
+T2A AudioX [43] 2.73 19.43 3.32 171.60 5.98 12.15 - - 0.28
+MMAudio [7] 2.54 11.25 5.07 142.71 5.54 9.28 - - 0.28
 […]
-V2A AudioX [tian2025audiox] 3.13 18.90 4.01 205.48 5.87 8.31 1.20 0.23 0.13
-MMAudio [cheng2025mmaudio] 2.94 13.41 3.87 159.30 5.50 7.59 0.62 0.24 0.14
+V2A AudioX [43] 3.13 18.90 4.01 205.48 5.87 8.31 1.20 0.23 0.13
+MMAudio [7] 2.94 13.41 3.87 159.30 5.50 7.59 0.62 0.24 0.14
 […]
-VT2A ThinkSound ($w/o$. CoT) [liu2025thinksound] 2.53 11.99 3.52 206.93 5.77 6.09 0.66 0.22 0.19
-HunyuanVideo-Foley [chen2024hunyuanfoley] 2.13 8.06 3.58 94.64 6.04 8.17 0.55 0.34 0.23
-AudioX [tian2025audiox] 2.39 14.26 3.16 149.37 5.97 10.23 1.21 0.23 0.26
-MMAudio [cheng2025mmaudio] 2.41 10.12 4.90 129.21 5.53 7.46 0.59 0.25 0.20
+VT2A ThinkSound ($w/o$. CoT) [32] 2.53 11.99 3.52 206.93 5.77 6.09 0.66 0.22 0.19
+HunyuanVideo-Foley [40] 2.13 8.06 3.58 94.64 6.04 8.17 0.55 0.34 0.23
+AudioX [43] 2.39 14.26 3.16 149.37 5.97 10.23 1.21 0.23 0.26
+MMAudio [7] 2.41 10.12 4.90 129.21 5.53 7.46 0.59 0.25 0.20
 […]
@@
 Datasets.
-For T2A backbone pre-training, we use a large-scale corpus comprising the train set of audio datasets such as AudioCaps [kim2019audiocaps], WavCaps [mei2023wavcaps], Clotho [drossos2020clotho], AudioSet [gemmeke2017audio], VGGSound [chen2020vggsound], FSD50k [fonseca2022fsd50k], as well as music datasets including MSD [bertin2011million] and FMA [defferrard2016fma].
+For T2A backbone pre-training, we use a large-scale corpus comprising the train set of audio datasets such as AudioCaps [24], WavCaps [36], Clotho [10], AudioSet [15], VGGSound [4], FSD50k [14], as well as music datasets including MSD [3] and FMA [9].
 To maintain consistency, all audio is segmented into 10-second clips and resampled at 16 kHz.
@@
 More details of the implementation are provided in Appendix Section G.
-For evaluation, we compare Omni2Sound with SOTA models on three benchmarks: our proposed VGGSound-Omni (Section 5), Kling-Audio-Eval [wang2025audiogenomni] and AudioCaps test set [kim2019audiocaps].
+For evaluation, we compare Omni2Sound with SOTA models on three benchmarks: our proposed VGGSound-Omni (Section 5), Kling-Audio-Eval [46] and AudioCaps test set [24].
 We strictly ensure that these evaluation benchmarks are strictly disjoint from all data used in our training stages to prevent potential data leakage.
 Evaluation Metrics.
-We implement our objective evaluation using the standardized AV-benchmark toolkit [hkchengrex2024avbenchmark] on 8-second clips, following previous work [cheng2025mmaudio].
-We assess quality across four critical dimensions [liu2023audioldm].
-For Distribution Matching, we measure feature similarity between generated and ground-truth audio using Fréchet Distance (FAD [hershey2017cnn], $\mathrm{FD}_{\mathrm{PaSST}}$ [koutini2022passt], FD [kong2020panns]) and Kullback-Leibler divergence (KL, $\mathrm{KL}_{\mathrm{PaSST}}$).
-Audio Quality is assessed via Inception Scores (IS [salimans2016improved], $\mathrm{IS}_{\mathrm{PaSST}}$) and Production Quality (PQ [tjandra2025meta]) for aesthetics.
-Semantic Alignment evaluates text-audio consistency (CLAP [elizalde2023clap], MS-CLAP [wu2023large]) and video-audio alignment (IB [girdhar2023imagebind]).
-Finally, Temporal Alignment is measured using the Desynchronization Score (DS) predicted by Synchformer [iashin2024synchformer].
+We implement our objective evaluation using the standardized AV-benchmark toolkit [7] on 8-second clips, following previous work [7].
+We assess quality across four critical dimensions [30].
+For Distribution Matching, we measure feature similarity between generated and ground-truth audio using Fréchet Distance (FAD [19], $\mathrm{FD}_{\mathrm{PaSST}}$ [26], FD [25]) and Kullback-Leibler divergence (KL, $\mathrm{KL}_{\mathrm{PaSST}}$).
+Audio Quality is assessed via Inception Scores (IS [39], $\mathrm{IS}_{\mathrm{PaSST}}$) and Production Quality (PQ [44]) for aesthetics.
+Semantic Alignment evaluates text-audio consistency (CLAP [11], MS-CLAP [48]) and video-audio alignment (IB [17]).
+Finally, Temporal Alignment is measured using the Desynchronization Score (DS) predicted by Synchformer [22].
 Detailed metric definitions and calculations are provided in the Appendix.
@@
 We present our main results on VGGSound-Omni benchmark in Table 3.
-To ensure a fair comparison, all baseline models are re-evaluated using their official checkpoints and the standardized AV-benchmark toolkit [hkchengrex2024avbenchmark], using the same video and text conditions.
-The results demonstrate that Omni2Sound achieves state-of-the-art performance across all three unified tasks (T2A, V2A, and VT2A) compared to both previous unified VT2A models (AudioX [tian2025audiox], MMAudio [cheng2025mmaudio]) and specialized models (e.g.
-ThinkSound [liu2025thinksound], HunyuanVideo-Foley [chen2024hunyuanfoley]).
-To further validate Omni2Sound’s generalization beyond our SoundAtlas captioning style, we evaluate it on the same VGGSound test clips but use the Video-LLaMA [zhang2023video-llama] captions from ThinkSound [liu2025thinksound].
+To ensure a fair comparison, all baseline models are re-evaluated using their official checkpoints and the standardized AV-benchmark toolkit [7], using the same video and text conditions.
+The results demonstrate that Omni2Sound achieves state-of-the-art performance across all three unified tasks (T2A, V2A, and VT2A) compared to both previous unified VT2A models (AudioX [43], MMAudio [7]) and specialized models (e.g.
+ThinkSound [32], HunyuanVideo-Foley [40]).
+To further validate Omni2Sound’s generalization beyond our SoundAtlas captioning style, we evaluate it on the same VGGSound test clips but use the Video-LLaMA [52] captions from ThinkSound [32].
 As shown in Table 2 (w/ Video-LLaMA caps), while performance sees a slight degradation, our model’s scores still surpass all baselines, confirming its robustness to unseen captioning styles.
 Generalization on Third-Party Benchmarks.
-To validate generalization, we evaluate on Kling-Audio-Eval [wang2025audiogenomni] and AudioCaps [kim2019audiocaps] results in Table 4 and Appendix Table 7.
+To validate generalization, we evaluate on Kling-Audio-Eval [46] and AudioCaps [24] results in Table 4 and Appendix Table 7.
 On Kling-Audio-Eval, Omni2Sound remains highly competitive despite the domain gap (YouTube-sourced SoundAtlas vs.
 Kling’s professional video).
-While trailing HunyuanVideo-Foley [chen2024hunyuanfoley] in some metrics, which is expected given its massive data advantage (100k vs 2k hours), our model consistently outperforms other unified and specialized baselines across all tasks.
+While trailing HunyuanVideo-Foley [40] in some metrics, which is expected given its massive data advantage (100k vs 2k hours), our model consistently outperforms other unified and specialized baselines across all tasks.
 Furthermore, on AudioCaps, Omni2Sound achieves top-tier performance against specialized T2A models, securing the best scores in distribution metrics ($\mathrm{KL}$, $\mathrm{FD}$) and semantic alignment ($\mathrm{CLAP}=0.36$), while remaining highly competitive in audio quality (PQ) and the $\mathrm{FAD}$ metric.
@@
 Our experiments demonstrate that this approach effectively resolves inter-task and intra-task competition and enables Omni2Sound to achieve unified state-of-the-art performance.
-[…]
 Supplementary Material
@@
 Appendix A Cost Analysis on Audio Captioning
-While Gemini 2.5 Pro [team2023gemini] represents a milestone as a native multimodal foundation model, utilizing it directly for large-scale video-grounded audio captioning proves economically unsustainable.
+While Gemini 2.5 Pro [42] represents a milestone as a native multimodal foundation model, utilizing it directly for large-scale video-grounded audio captioning proves economically unsustainable.
 As quantified in Table 6, using Gemini’s standard API pricing, a naive implementation—processing raw video frames alongside audio (V+AV+A)—incurs a prohibitive expenditure of \$10,275 USD per 1M samples.
@@
 Method KL FD FAD PQ LA-CLAP
-AudioLDM 2-L [liu2024audioldm2] 1.73 34.21 2.26 5.93 0.24
-TANGO 2 [majumder2024tango2] 1.19 15.92 3.17 5.82 0.35
-Make-An-Audio 2 [huang2023makeanaudio2_placeholder] 1.38 15.34 1.46 5.64 0.25
-GenAU-Large [haji-ali2024genau] 1.42 16.92 1.32 5.52 0.26
-MMAudio [cheng2025mmaudio] 1.43 13.78 2.92 5.30 0.29
-AudioX [tian2025audiox] 1.55 17.10 2.65 5.81 0.31
+AudioLDM 2-L [31] 1.73 34.21 2.26 5.93 0.24
+TANGO 2 [35] 1.19 15.92 3.17 5.82 0.35
+Make-An-Audio 2 [20] 1.38 15.34 1.46 5.64 0.25
+GenAU-Large [18] 1.42 16.92 1.32 5.52 0.26
+MMAudio [7] 1.43 13.78 2.92 5.30 0.29
+AudioX [43] 1.55 17.10 2.65 5.81 0.31
 […]
@@
 Synthetic Music Augmentation.
-To address the high demand for Background Music (BGM) generation, we create a Synthetic Music subset by mixing semantically aligned MusicCaps [agostinelli2023musiccaps] clips into a pool of high-fidelity videos.
+To address the high demand for Background Music (BGM) generation, we create a Synthetic Music subset by mixing semantically aligned MusicCaps [1] clips into a pool of high-fidelity videos.
 This process follows a two-stage procedure.
@@
 Comparison with Concurrent Work.
-We acknowledge the pioneering work of VinTAGe-Bench [bain2025vintage] in synthetic robustness evaluation.
+We acknowledge the pioneering work of VinTAGe-Bench [29] in synthetic robustness evaluation.
 However, the off-screen subset of our VGGSound-Omni benchmark extends this direction in three critical dimensions.
-First, in terms of Realism, by leveraging VGGSounder [zverev2025vggsounder] metadata, our natural subset is primarily sourced from real-world off-screen audio events rather than relying solely on synthetic mixes.
+First, in terms of Realism, by leveraging VGGSounder [54] metadata, our natural subset is primarily sourced from real-world off-screen audio events rather than relying solely on synthetic mixes.
 Second, regarding Scale, our benchmark is significantly larger, providing 1,613 evaluation items compared to the 212 basic videos of VinTAGe-Bench.
@@
 Appendix E Generalization on Third-Party Benchmarks.
-To further validate our model’s generalization and mitigate potential biases from our self-constructed benchmark, we evaluate it on the Kling-Audio-Eval [wang2025audiogenomni] and Audiocaps test set [kim2019audiocaps].
+To further validate our model’s generalization and mitigate potential biases from our self-constructed benchmark, we evaluate it on the Kling-Audio-Eval [46] and Audiocaps test set [24].
 In Table 4, on the Kling-Audio-Eval benchmark, Omni2Sound remains highly competitive, despite a significant data scale and distribution gap (our YouTube-sourced SoundAtlas vs.
 Kling’s professional video/Foley).
-While HunyuanVideo-Foley [chen2024hunyuanfoley] leads on several metrics, this is expected given its massive 100k-hour internal dataset, which is tens of times larger than our SoundAtlas filter derived from VGGSound and AudioSet.
+While HunyuanVideo-Foley [40] leads on several metrics, this is expected given its massive 100k-hour internal dataset, which is tens of times larger than our SoundAtlas filter derived from VGGSound and AudioSet.
 Nevertheless, Omni2Sound consistently outperforms all other strong baselines (e.g., MMAudio, AudioX, and ThinkSound) across V2A and VT2A tasks, demonstrating strong generalization as the SOTA or second-best method.
@@
 Model Configuration.
-Following Stable Audio [evans2024stable], our diffusion model adopts a Diffusion Transformer (DiT) architecture within a Latent Diffusion Model (LDM) paradigm.
+Following Stable Audio [13], our diffusion model adopts a Diffusion Transformer (DiT) architecture within a Latent Diffusion Model (LDM) paradigm.
 The diffusion backbone consists of a DiT with 24 layers, 24 attention heads, and a hidden dimension of 1536.
@@
 Both the conditional token dimension and the global condition embedding dimension are 1024.
-Finally, for audio compression, we train a Variational Autoencoder (VAE) from scratch based on the wav Audio VAE architecture [evans2024stable], operating at a 16kHz sampling rate.
+Finally, for audio compression, we train a Variational Autoencoder (VAE) from scratch based on the wav Audio VAE architecture [13], operating at a 16kHz sampling rate.
 With strides of $[4,4,4,10]$, the encoder achieves a total downsampling ratio of 640, mapping mono waveforms into a compact 64-dimensional latent space.
@@
 Training Data.
-For T2A backbone pre-training, we use a large-scale corpus comprising the train set of audio datasets such as AudioCaps [kim2019audiocaps], WavCaps [mei2023wavcaps], Clotho [drossos2020clotho], AudioSet [gemmeke2017audio], VGGSound [chen2020vggsound], FSD50k [fonseca2022fsd50k], as well as music datasets including MSD [bertin2011million] and FMA [defferrard2016fma].
+For T2A backbone pre-training, we use a large-scale corpus comprising the train set of audio datasets such as AudioCaps [24], WavCaps [36], Clotho [10], AudioSet [15], VGGSound [4], FSD50k [14], as well as music datasets including MSD [3] and FMA [9].
 All audio signals are standardized to a mono-channel format at 16kHz.
@@
 Appendix H Objective Evaluation Metrics.
-We implement our objective evaluation metrics using the standardized AV-benchmark toolkit [hkchengrex2024avbenchmark].
-All samples are generated under the same video and text conditions and evaluated in 8-second clips, following previous work [cheng2025mmaudio].
-Following common practice [liu2023audioldm], we assess the quality of the generation in four critical dimensions.
+We implement our objective evaluation metrics using the standardized AV-benchmark toolkit [7].
+All samples are generated under the same video and text conditions and evaluated in 8-second clips, following previous work [7].
+Following common practice [30], we assess the quality of the generation in four critical dimensions.
 For Distribution Matching, we measure the similarity in feature distribution between generated and ground-truth audio.
-We compute the Fréchet Distance using the VGGish (FAD) [hershey2017cnn] and PaSST ($\mathrm{FD}_{\mathrm{PaSST}}$) [koutini2022passt] embeddings, as well as the Fréchet Audio Distance using PANNs (FD) [kong2020panns].
+We compute the Fréchet Distance using the VGGish (FAD) [19] and PaSST ($\mathrm{FD}_{\mathrm{PaSST}}$) [26] embeddings, as well as the Fréchet Audio Distance using PANNs (FD) [25].
 We also report the Kullback-Leibler divergence using PANNs (KL) and PaSST ($\mathrm{KL}_{\mathrm{PaSST}}$) classifiers.
-For Audio Quality, we assess the quality of the generation using the Inception Score [salimans2016improved], calculated with both the PANNs (IS) and PaSST ($\mathrm{IS}_{\mathrm{PaSST}}$) classifiers.
-For Semantic Alignment, we evaluate text-audio consistency using LAION CLAP (CLAP) [elizalde2023clap] and Microsoft CLAP (MS-CLAP) [wu2023large] scores, and video-audio alignment using ImageBind score (IB) [girdhar2023imagebind] as cosine similarity between video and audio embeddings.
-Finally, for Temporal Alignment, we assess audio-visual synchrony using the DS metric predicted by Synchformer [iashin2024synchformer].
+For Audio Quality, we assess the quality of the generation using the Inception Score [39], calculated with both the PANNs (IS) and PaSST ($\mathrm{IS}_{\mathrm{PaSST}}$) classifiers.
+For Semantic Alignment, we evaluate text-audio consistency using LAION CLAP (CLAP) [11] and Microsoft CLAP (MS-CLAP) [48] scores, and video-audio alignment using ImageBind score (IB) [17] as cosine similarity between video and audio embeddings.
+Finally, for Temporal Alignment, we assess audio-visual synchrony using the DS metric predicted by Synchformer [22].
 Audio Captioning Instruction for SoundAtlas Roles and Tasks You are an experienced audio content analyst skilled in describing soundscapes through detailed, multi-dimensional natural language.